# Theoretical Modeling of LLM Self-Improvement Training Dynamics Through Solver-Verifier Gap

### v1

@@ 5.2 Experiments
Table 1: Solver accuracy with QE method on train set: raw data and relative improvements of strategies average (%). We note that the improvement of late strategy on Math dataset is slight, which may caused by the insufficient training on the external data.
Strategy Phi-4-mini-instruct Llama-3.2-3B-Instruct
Math GSM8k Math GSM8k
@@
Baseline 43.87 ($\pm$ 0.36) 87.71 ($\pm$ 0.11) 49.16 ($\pm$ 0.73) 87.00 ($\pm$ 0.21)
Early 44.59 ($\pm$ 0.86) 87.72 ($\pm$ 0.24) 52.52 ($\pm$ 0.47) 86.35 ($\pm$ 0.32)
Uniform 45.21 ($\pm$ 0.50) 87.82 ($\pm$ 0.21) 52.73 ($\pm$ 0.34) 85.99 ($\pm$ 0.65)
Late 43.73 ($\pm$ 0.13) 87.71 ($\pm$ 0.21) 48.98 ($\pm$ 0.17) 85.07 ($\pm$ 0.33)
Max-Min 1.48 0.11 3.75 1.28
Avg 44.51 87.75 51.41 85.80
Avg vs Initial +14.20 +14.33 +15.39 +22.70
Avg vs Baseline +0.64 +0.04 +2.25 -1.20
Early($t=0$) vs Initial (Verifier) +5.87 +0.96 +0.97 -1.62
In this section, we perform cross-improvement experiments using different allocation strategies.  Figure 7 illustrates the process of cross-improvement. For these experiments, external data are generated by DeepSeek-V3. Specifically, we utilize DeepSeek-V3 responses for fine-tuning instead of BoN responses. We perform $10$ epochs of cross-improvement on the Phi4-mini and Llama3.2-3B models using the QE method. To facilitate the model’s learning from external data, all such data are introduced within the first $8$ training epochs; data introduced in an epoch remain available throughout $10$-epoch period. The total number of DeepSeek-V3 responses used is $3000$, and we test three allocation strategies:
@@
  • Early: All external data are introduced in the first epoch.
  • Uniform: An equal amount of new external data is introduced in each of the first eight epochs.
  • Late: All external data are introduced in the eighth epoch.
We present the results on the training data in Table 1. In this table, Initial represents the performance of the original model, while Baseline represents the results from self-improvement. We observe that the average solver accuracy for all three strategies is higher than the baseline on the MATH dataset. However, Phi4-mini shows a slight performance improvement on GSM8k, while Llama3.2-3B even exhibits a performance drop on this dataset.
To better understand why cross-improvement is less effective on GSM8k, we calculate the difference in verifier accuracy at $t=0$ between the Early strategy and the Initial model, where $t=0$ signifies that external data are incorporated but no training is performed. This difference is listed in the last row of Table 1. We find that this difference is small for Phi4-mini and negative for Llama3.2-3B. This implies that, on GSM8k, the verifier’s performance after the addition of external data is not significantly better than its performance without such data, and may even decrease. This observation indicates that a key factor influencing the effectiveness of cross-improvement is indeed whether the external model’s verifier capability significantly surpasses that of the original model.
Additionally, we also observe that the late strategy has only a slight improvement on the Math dataset. To determine the cause of this phenomenon, we conduct a $12$ epochs cross-improvement experiment on the Phi4-mini and Llama3.2-3B model and present the result in Table 2. The solver accuracy of late strategy rises to $45.83\%$ and $51.31\%$, respectively, indicating that the model has not fully learned the external data with late strategy in the original experiment. Furthermore, we observe that the performance differences in the solver accuracy between the three strategies are marginal. This suggests that the timing of external data integration does not significantly influence the effectiveness of the cross-improvement method.
Table 2: Solver accuracy of three strategies with QE method on train set for $12$ epoch-training. The result demonstrates that late strategy has similar result with early and uniform strategy after $12$-epoch training.
Strategy Phi-4-mini-instruct Llama-3.2-3B-Instruct
Initial 30.31 ($\pm$ 0.24) 36.02 ($\pm$ 0.25)
Baseline 43.87 ($\pm$ 0.36) 49.16 ($\pm$ 0.73)
Early 46.33 ($\pm$ 0.13) 53.47 ($\pm$ 0.33)
Uniform 46.56 ($\pm$ 0.15) 52.73 ($\pm$ 0.66)
Late 45.83 ($\pm$ 0.39) 51.31 ($\pm$ 0.62)
6 Conclusion
In this paper, we propose a theoretical framework to analyze the training dynamics of self-improvement via the solver-verifier gap. Experimental results on various datasets and models accord well with the theoretical findings. Besides, one may derive the theoretical limits based on the framework, which is closely related to the model’s verification capability. To break the limit, one may apply cross-improvement to enhance the model’s verification capability. Therefore, we further introduce the corresponding theoretical framework on the dynamics of cross-improvement under limited external data regimes, and find that the allocation of external data might have less influence on the final results. Experimental results verify the theoretical findings. In the end, we provide several potential future directions:
  • Although the experiments accord well with the theory, the mechanism behind the theoretical framework is still under-explored;
  • We assume a time-invariant property when analyzing the cross-improvement, which might be relaxed in future work.
  • Developing self-improvement algorithms based on our framework is a potential direction in the future.
  • External data can be used to fine-tune a well-self-improved model to further improve the performance of the model.
Appendix
This section provides the supplementary material. Appendix A provides experimental details and omitted experimental results. Appendix B provides omitted derivations of the theoretical models. Appendix C discuss the difference between our work and prior works, as well as the challenges and future direction for self-improvement.
Appendix A Experimental Details
In this section, we detail our experiment hyperparameters and present all results of our experiment. In Section A.1, we detail the hyperparameters chosen in our experiments. We provide all All of our experiments are run on 80G NVIDIA A800 GPUs.
A.1 Hyperparameters
In this section, we detail the hyperparameters in Table 3.
Table 3: Hyperparameters for SFT
Learning Rate Weight Decay LoRA Rank LoRA dropout Solver Temperature Verifier Temperature
1e-5 0.01 16 0.5 1 0.1
A.2 Omitted figures
In this section, we will provide figures omitted in Section 4.
Self-improvement In this part, we display figures omitted in Section 4Figure 8 illustrates the results of self-improvement on Phi3.5-mini, Phi3-mini and Llama3.2-3B model with QE method. We observe that most of model-task pairs have similar results with the Phi4-mini model except for several pairs. We also present the results with TF method in Figure 9. When self-improving the Phi3.5-mini and Phi3-mini models on Math data set, we observe that accuracy and uncertainty decrease simultaneously after several training epochs. The reason for this phenomenon might be LLM misleading by incorrect responses, as the BoN response may correspond to a incorrect answer with low uncertainty. The solution of this problem could be a future direction. Additionally, we note that the verifier accuracy of Llama3.2-3B on GSM8k decreases as training progresses and is lower than the solver accuracy after 8 epochs of training. One possible reason for this phenomenon is that we use length-regularized log-likelihood to obtain the BoN responses, so the BoN responses tend to be longer responses. Long responses are more likely to contain repeated content, which may reduce accuracy.
Self-evaluation This part we provide the results of self-evaluation for different sample size $N$Figure 10 illustrates the accuracy and uncertainty of Phi3.5-mini, Phi3-mini, Llama3.2-3B and Llama3.1-8B on Math and GSM8k with different sample size using TF and QE respectively, which shows that the verifier outperforms solver in all model-task pairs.
Cross-evaluation We present cross-evaluation results in Figure 11. It shows that the verifier always performs better than solver, although they use different LLMs.
A.3 Curve fitting
In this section, we present the fitted curves of Phi3.5-mini, Phi3-mini and Llama3.2-3B in Figure 12 and observe that all curves fitted well.
A.4 Pass@K
In this section, we present the Pass@K with TF method at different training epochs in Figure 13. We observe that Pass@K increases when K is small and decreases when K is large, which is similar to the results with the QE method.
(a) Phi3.5-mini
(b) Phi3-mini
(c) Llama3.2-3B
Figure 8: Accuracy and uncertainty during the self-improvement of the Phi3.5-mini, Phi3-mini and Llama3.2-3B on the Math and GSM8k datasets using the QE method.
(a) Phi4-mini
(b) Phi3.5-mini
(c) Phi3-mini
(d) Llama3.2-3B
Figure 9: Accuracy and uncertainty during the self-improvement of the Phi4-mini, Phi3.5-mini, Phi3-mini and Llama3.2-3B on the Math and GSM8k datasets using the TF method.
(a) Phi3.5-mini
(b) Phi3-mini
(c) Llama3.2-3B
(d) Llama3.1-8B
Figure 10: Accuracy and uncertainty of Phi3.5-mini, Phi3-mini, Llama3.2-3B and Llama3.1-8B on Math and GSM8k with different sample size using TF and QE respectively.
Figure 11: (a) Cross-evaluation using TF method with sample size N=16 on Math (b) Cross-evaluation using QE method with sample size N=16 on GSM8k (c) Cross-evaluation using TF method with sample size N=16 on GSM8k. For each solver model, sampled responses are verified by different models. Figures on the left illustrate the difference of accuracy while figures on the right show the 10 logarithms of the uncertainty difference.
(a) Phi3.5-mini
(b) Phi3-mini
(c) Llama3.2-3B
Figure 12: Uncertainty curve versus epochs of Phi4-mini, Phi3.5-mini, Phi3-mini and Llama3.2-3B models using QE and TF methods on Math and GSM8k datasets.
Figure 13: Pass@K with TF method at $t=0$, $t=5$ and $t=10$.
@@ Appendix B Theoretical Derivation
In this section, we present the theoretical derivations of the conclusion presented in Section 5.1, from the dynamics conditions. The discussion and conclusion presented at the end of Section 5.1 stay the same.
According to Section 5.1, the dynamics conditions are formulated as:
@@
With these notations, we can derive the following iteration from Equation 22, 23 and 24:
@@
where $\mathbf{U}(0)$ denotes $\begin{bmatrix}U_{v,0},&U_{s,0},&1\end{bmatrix}^{\top}$. The meaning of $U_{v,0},U_{s,0}$ is defined in Equation 24.
Under the circumstances, the following approximation holds true:
$\prod_{t=1}^{T}(I-\mathbf{\Delta}_{t})\approx\prod_{t=1}^{T}e^{-\mathbf{\Delta}_{t}}\approx e^{-\sum_{t=1}^{T}\mathbf{\Delta}_{t}}\approx e^{-\mathbf{\Delta^{\prime}}},$ (29)
@@
The third approximation is based on the approximation $\frac{1}{1+\gamma\eta_{t}}\approx 1-\gamma\eta_{t}$, given $\eta_{t}$ a small quantity.
@@
Equation 29 and 30 indicates that one can derive a solution of $U_{s}(T)$, which is only related to $\sum_{t=1}^{T}\eta_{t}$. Therefore, one can come to the conclusion that (i)solver uncertainty at the final epoch $T$ is approximately the same for cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, and (ii)with the two elements in matrix $-\mathbf{\Delta^{\prime}}$ that $\gamma$ appears in both negatively correlated to the term $\gamma\sum_{t=1}^{T}\eta_{t}$, the final solver uncertainty is also negatively correlated to this term. This indicates that cross-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=0$, in terms of solver capability.
Appendix C Review and Prospect
Future: (i) We propose a new framework for self-improvement, which provides a new perspective on self-improvement. Developing self-improvement algorithms based on our framework is a potential direction in the future. (ii) As there are no significant differences in the results of different allocation strategies, external data can be used to fine-tune a well-self-improved model to further improve the performance of the model. Compared with cross-improvement of the initial model, this method could reduce the resource consumption of training and obtain similar results.

### v3

@@ 5.2 Experiments
Table 1: Solver accuracy with QE method on train split: raw data and relative improvements of strategies average (%). Initial represents the performance of the original model, while Baseline represents the results from self-improvement. We deploy three strategies: Early, Uniform, and Late. The last row shows the difference in verifier accuracy between using all external data at the start (Early strategy at $t=0$) and the baseline’s initial verifier accuracy (Initial).
Strategy Phi-4-mini Llama-3.2-3B
Math(%) GSM8k(%) Math(%) GSM8k(%)
@@
Baseline 45.08 ($\pm$ 0.48) 88.53 ($\pm$ 0.48) 49.16 ($\pm$ 0.90) 88.66 ($\pm$ 0.73)
Early 46.33 ($\pm$ 0.16) 88.54 ($\pm$ 0.12) 53.48 ($\pm$ 0.41) 88.26 ($\pm$ 0.57)
Uniform 46.56 ($\pm$ 0.19) 88.44 ($\pm$ 0.05) 52.74 ($\pm$ 0.82) 88.23 ($\pm$ 0.58)
Late 45.83 ($\pm$ 0.48) 88.90 ($\pm$ 0.21) 51.31 ($\pm$ 0.76) 87.19 ($\pm$ 0.28)
Max-Min 0.73 0.46 2.17 1.07
Avg 46.24 88.63 52.51 87.89
Avg vs Initial +15.93 +15.21 +16.49 +24.79
Avg vs Baseline +1.16 +0.10 +3.35 -0.77
Early($t=0$) vs Initial (Verifier) +5.87 +0.96 +0.97 -1.58
In this section, we perform cross-improvement experiments using different allocation strategies.  Figure 6 illustrates the process of cross-improvement. For these experiments, external data are generated by DeepSeek-V3 which achieves accuracy of $60.47\%$ and $91.18\%$ on Math and GSM8k dataset respectively. Specifically, we utilize DeepSeek-V3 responses for fine-tuning instead of BoN responses. We test three allocation strategies: (a) Early: All external data are introduced in the first epoch. (b) Uniform: An equal amount of new external data is introduced in each epoch. (c) Late: All external data are introduced in the eighth epoch.
We present the results on the training data in Table 1. We observe that the average solver accuracy for all three strategies is higher than the baseline on the MATH dataset. However, Phi-4-mini shows a slight performance improvement on GSM8k, while Llama-3.2-3B exhibits a drop on GSM8k dataset. This observation indicates that a key factor influencing the effectiveness of cross-improvement is indeed whether the external model’s verifier capability significantly surpasses that of the original model. In addition, we find that the difference in accuracy between the three strategies is slight, which validates the conclusions under proposition 5.1.
6 Conclusion
In this paper, we propose a theoretical framework to analyze the training dynamics of self-improvement via the solver-verifier gap. Experimental results on various datasets and models accord well with the theoretical findings. Besides, one may derive the theoretical limits based on the framework, which is closely related to the model’s verification capability. To break the limit, one may apply cross-improvement to enhance the model’s verification capability. Therefore, we further introduce the corresponding theoretical framework on the dynamics of cross-improvement under limited external data regimes, and find that the allocation of external data might have less influence on the final results. Experimental results verify the theoretical findings.
In the end, we provide several future directions: (i) We assume a time-invariant property when analyzing the cross-improvement, which might be relaxed in future work; (ii) External data can be used to fine-tune a self-improved model to further improve the performance of the model; (iii) We use a phenomenological modeling approach in our framework. A deeper investigation into the underlying mechanisms that cause the self-improvement dynamics to be well-described by our physics-inspired model would be a key future direction. This includes further research into the relationship between the model’s key parameters, $\alpha$ and $\beta$, and factors such as the LLM’s architecture.
Appendix
This part provides the supplementary materials. Appendix A provides necessary details omitted before. Appendix B provides omitted derivations of the theoretical framework. Appendix C provides experiment details and omitted experiment results.
Appendix A Supplementary Details
In this section, we provide the details omitted before. Section A.1 presents the omitted illustration. We discuss the difference between our work and prior works, as well as the challenges for self-improvement in Section A.2.
A.1 Omitted illustration
Figure 5 gives an overview of our theoretical framework while Figure 6 illustrates cross-improvement process.
Figure 5: An overview of our theoretical framework.
@@ A.1 Omitted illustration
A.2 Review and Prospect
@@ Appendix B Theoretical Derivation
In this section, we present the detailed theoretical derivations omitted before.
B.1 Proof of proposition 3.1
In this section, we present the detailed derivation of proposition 3.1. We first derive Equation 11. From Equation 9 and linear approximation $f(G)\approx kG-b$, we have
$\frac{dG(t)}{dt}\approx-(\alpha-\beta)(kG-b).$ (19)
The general solution of Equation 19 is
$G(t)\approx Ce^{-k(\alpha-\beta)t}+\frac{b}{k},$ (20)
where C is a constant. To find the constant $C$, we use the initial condition at $t=0$: $G(0)=U_{s,0}-U_{v,0}=G_{0}$, thus
$G(0)=Ce^{0}+\frac{b}{k},$ (21)
$C=U_{v,0}-U_{s,0}-\frac{b}{k}.$ (22)
Let $\delta=U_{v,0}-U_{s,0}-\frac{b}{k}$, we have
$G(t)\approx\delta e^{-k(\alpha-\beta)t}+\frac{b}{k}.$ (23)
Then we derive Equation 12. From Equation 23, we have
$E(t)\approx kG(t)-b=k(\delta e^{-k(\alpha-\beta)t}+b/k)-b=k\delta e^{-k(\alpha-\beta)t}.$ (24)
Solving the original differential equation for $U_{s}(t)$:
$\frac{dU_{s}(t)}{dt}=-\alpha E(t)\approx-\alpha(k\delta e^{-k(\alpha-\beta)t}),$ (25)
$\int dU_{s}(t)=\int-\alpha k\delta e^{-k(\alpha-\beta)t}dt,$ (26)
$\int-\alpha k\delta e^{-k(\alpha-\beta)t}dt=-\alpha k\delta\left(\frac{e^{-k(\alpha-\beta)t}}{-k(\alpha-\beta)}\right)+C_{s},$ (27)
$U_{s}(t)\approx\frac{\alpha\delta}{\alpha-\beta}e^{-k(\alpha-\beta)t}+C_{s}.$ (28)
Using the initial condition $U_{s}(0)=U_{s,0}$ to solve for the constant $C_{s}$:
$U_{s,0}=\frac{\alpha\delta}{\alpha-\beta}e^{0}+C_{s}\implies C_{s}=U_{s,0}-\frac{\alpha\delta}{\alpha-\beta},$ (29)
Let $\alpha^{\prime}=\frac{\alpha\delta}{\alpha-\beta}$ and $U_{s,\infty}=U_{s,0}-\alpha^{\prime}$. Thus, the solution for $U_{s}(t)$ is
$U_{s}(t)\approx\alpha^{\prime}e^{-k(\alpha-\beta)t}+U_{s,\infty}.$ (30)
The derivation for $U_{v}(t)$ is analogous, yielding
$U_{v}(t)\approx\beta^{\prime}e^{-k(\alpha-\beta)t}+U_{v,\infty},$ (31)
where $\beta^{\prime}=\frac{\beta\delta}{\alpha-\beta}$ and $U_{v,\infty}=U_{v,0}-\beta^{\prime}$.
B.2 Proof of corollary 3.1
In this section, we present the detailed derivation of corollary 3.1. From proposition 3.1, the final solver uncertainty is given by $U_{s,\infty}=U_{s,0}-\alpha^{\prime}$, where $\alpha^{\prime}=\frac{\alpha}{\alpha-\beta}\delta$ and $\delta=G_{0}-b/k$. Substituting these into the expression for $U_{s,\infty}$ yields $U_{s,\infty}=U_{s,0}-\frac{\alpha}{\alpha-\beta}(G_{0}-\frac{b}{k})$. Assuming $U_{s,0}$ is constant, we have $\frac{\partial U_{s,0}}{\partial G_{0}}=0$. Thus, we have $\frac{\partial U_{s,\infty}}{\partial G_{0}}=-\frac{\alpha}{\alpha-\beta}$.
B.3 Proof of corollary 3.2
In this section, we present the detailed derivation of corollary 3.2. The objective of corollary 3.2 is to find the time $t$ such that the capability gap $G(t)$ is within a tolerance $\epsilon$ of its convergence value $G_{\infty}$, which is expressed as:
$|G(t)-G_{\infty}|<\epsilon.$ (32)
Starting from the expression for $G(t)$ given in proposition 3.1, we have:
$G(t)\approx\delta e^{-k(\alpha-\beta)t}+G_{\infty},$ (33)
$G(t)-G_{\infty}\approx\delta e^{-k(\alpha-\beta)t}.$ (34)
Since $\delta$ is defined on the basis of the initial gap, we can assume $\delta>0$. Thus, we have
$\delta e^{-k(\alpha-\beta)t}<\epsilon,$ (35)
$t>\frac{\ln(\delta/\epsilon)}{k(\alpha-\beta)}.$ (36)
B.4 Proof of proposition 5.1
In this section, we present the theoretical derivation of proposition 5.1. According to Section 5.1, the dynamics conditions are formulated as:
@@
With these notations, we can derive the following iteration from Equation 38, 39 and 40:
@@
where $\mathbf{U}(0)$ denotes $\begin{bmatrix}U_{v,0},&U_{s,0},&1\end{bmatrix}^{\top}$. The meaning of $U_{v,0},U_{s,0}$ is defined in Equation 40. Under the circumstances, the following approximation holds true:
$\prod_{t=1}^{T}(I-\mathbf{\Delta}_{t})\approx\prod_{t=1}^{T}e^{-\mathbf{\Delta}_{t}}=e^{-\sum_{t=1}^{T}\mathbf{\Delta}_{t}}=e^{-\mathbf{\Delta^{\prime}}},$ (45)
@@
The third approximation is based on the approximation $\frac{1}{1+\gamma\eta_{t}}\approx 1-\gamma\eta_{t}$, given $\eta_{t}$ a small quantity. Equation 45 and 46 indicates that we can derive a solution of $U_{s}(T)$ only related to $\sum_{t=1}^{T}\eta_{t}$. Therefore, we come to the conclusion that (i) solver uncertainty at the final epoch $T$ is approximately the same for cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, and (ii) with the two elements in matrix $\mathbf{\Delta^{\prime}}$ that $\gamma$ appears in both negatively correlated to the term $\gamma\sum_{t=1}^{T}\eta_{t}$, the final solver uncertainty is also negatively correlated to this term. This indicates that cross-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=0$, in terms of solver capability.
Appendix C Experimental Details
In this section, we detail our experiment hyperparameters and present all results of our experiment. In Section C.1, we detail the hyperparameters chosen in our experiments. We provide all experiment results in Section C.2. All of our experiments are run on 80G NVIDIA A800 GPUs.
C.1 Hyperparameters
In this section, we detail the hyperparameters in Table 2.
Table 2: Hyperparameters for SFT
Learning Rate Weight Decay Solver Temperature Verifier Temperature
1e-5 0.01 1 0.1
Sample Size (N) Max New Tokens LoRA Rank LoRA dropout
16 512 16 0.5
C.2 Omitted Figures
In this section, we will provide experiment details including setup and figures omitted in Section 4.
Verification Methods. We use two verification methods TrueFalse and Quality Evaluation in the experiment. TrueFalse (TF): The solver generates $N$ responses, denoted $\hat{y}_{i,1},\cdots,\hat{y}_{i,N}$, for a prompt $x_{i}$. The verifier is then tasked with answering whether each response $\hat{y}_{i,j}$ is correct. If the verifier deems a response $\hat{y}_{i,j}$ correct, its score $s(\hat{y}_{i,j})$ is set to $1$; otherwise, it is set to 0; Quality Evaluation (QE): The solver generates $N$ responses, $\hat{y}_{i,1},\cdots,\hat{y}_{i,N}$, for a prompt $x_{i}$. The verifier then assigns a continuous score $s(\hat{y}_{i,j})$ between 0 and 1 to each response based on its quality. A score of $s(\hat{y}_{i,j})=0$ indicates a completely incorrect answer, while $s(\hat{y}_{i,j})=1$ indicates a completely correct answer.
Self Improvement. In this part, we display figures omitted in Section 4. In self-improvement, we employed mini-batch gradient descent with a batch size of $256$. For each model-task pair, we conducted training for $10$ epochs, saving a checkpoint after processing half of the training data. We test the solver accuracy and the verifier accuracy, defined as the accuracy of the response and the BoN response, respectively. Figure 8 illustrates the results of self-improvement on Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B model with QE method. We observe that most of model-task pairs have similar results with the Phi-4-mini model except for several pairs. We also present the results with TF method in Figure 9. When self-improving the Phi-3.5-mini and Phi-3-mini models on Math data set, we observe that accuracy and uncertainty decrease simultaneously after several training epochs. The reason for this phenomenon might be LLM misleading by incorrect responses, as the BoN response may correspond to a incorrect answer with low uncertainty. The solution of this problem could be a future direction. Additionally, we note that the verifier accuracy of Llama-3.2-3B on GSM8k decreases as training progresses and is lower than the solver accuracy after 8 epochs of training. One possible reason for this phenomenon is that we use length-regularized log-likelihood to obtain the BoN responses, so the BoN responses tend to be longer responses. Long responses are more likely to contain repeated content, which may reduce accuracy. In addition, we present the self-improvement results on ProntoQA dataset with TF method Figure 10.
Validation. In this part, we present the validation results of Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on train split in Figure 11 and observe that the empirical data are strongly fitted to the exponential model for all models. Results of Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on test split are presented in Figure 12. Experiment results on ProntoQA dataset are presented in Figure 13.
Self Evaluation. This part we provide the results of self-evaluation for different sample size $N$. For each model, we evaluate the accuracy and uncertainty of both the solver and the verifier, varying $N$ across the values 2,4,8,16,32,64{2,4,8,16,32,64}.  Figure 14 illustrates the accuracy and uncertainty of Phi-4-mini, Phi-3.5-mini, Phi-3-mini, Llama-3.2-3B and Llama-3.1-8B on Math and GSM8k with different sample size using TF and QE respectively, which shows that the verifier outperforms solver in all model-task pairs.
Cross Evaluation. In cross evaluation we set $N=16$ because, as demonstrated in Figure 14, the accuracy improve slightly for $N>16$. The results of QE method are presented in Figure 16 while results of TF method are presented in Figure 17. We observe that when a fixed model serves as the solver, the verifier’s performance generally surpasses that of the solver, even when a different model is employed as the verifier.
Pass@K. We investigate the underlying reason for the limit of self-improvement, focusing on how a decrease in the model’s response diversity leads to a diminishing potential energy, thereby causing the model’s capability to plateau. Although the self-improvement paradigm shows great improvement in model capability, it iteratively leads to a performance plateau. To investigate the underlying cause of this saturation, we conduct the Pass@K experiment on the initial model and the self-improved model. Pass@K is a metric that calculates the proportion of prompts where at least one of the $K$ responses is correct. When $K$ is large, Pass@K could be used to measure the diversity of the solver. We present the result in Figure 18. We observe that when $K$ is small, Pass@K increases with the number of epochs of self-improvement, validating the self-improvement process. However, when $K$ is large, we observe a slight decrease in Pass@K, indicating that the diversity of the solver is reduced through self-improvement. The degradation in diversity is caused by the convergence to a certain response, which is a potential reason for the limit of self-improvement. Lack of ProntoQA dataset in experiment as the Pass@K of the QA problem is close to $1$ when $K$ is large.
(a) Phi-3.5-mini
(b) Phi-3-mini
(c) Llama-3.2-3B
Figure 7: Validation of the linear relationship between the Uncertainty gap $G$ and its rate of change $dG/dt$ on Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B with QE method. The scatter points represent empirical data from self-improvement on the Math and GSM8k datasets, while the solid lines show the linear regression fits.
(a) Phi-3.5-mini
(b) Phi-3-mini
(c) Llama-3.2-3B
Figure 8: Accuracy and uncertainty during the self-improvement of the Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on the Math and GSM8k datasets using the QE method.
(a) Phi-4-mini
(b) Phi-3.5-mini
(c) Phi-3-mini
(d) Llama-3.2-3B
Figure 9: Accuracy and uncertainty during the self-improvement of the Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on the Math and GSM8k datasets using the TF method.
Figure 10: Accuracy and uncertainty during the self-improvement of the Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on ProntoQA datset using the TF method.
(a) Phi-3.5-mini
(b) Phi-3-mini
(c) Llama-3.2-3B
Figure 11: Uncertainty curve versus epochs of Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B models using QE and TF methods on Math and GSM8k datasets train split.
(a) Phi-4-mini
(b) Phi-3.5-mini
(c) Phi-3-mini
(d) Llama-3.2-3B
Figure 12: Uncertainty curve versus epochs of Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B models using QE and TF methods on Math and GSM8k datasets test split.
(a) Train Split
(b) Test Split
Figure 13: Uncertainty curve versus epochs of Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B models using TF method on ProntoQA datasets.
(a) Phi-4-mini
(b) Phi-3.5-mini
(c) Phi-3-mini
(d) Llama-3.2-3B
(e) Llama-3.1-8B
Figure 14: Accuracy and uncertainty of Phi-4-mini, Phi-3.5-mini, Phi-3-mini, Llama-3.2-3B and Llama-3.1-8B on Math and GSM8k across different sample size using TF and QE respectively.
Figure 15: Accuracy and uncertainty of Phi-4-mini, Phi-3.5-mini, Phi-3-mini and Llama-3.2-3B on ProntoQA dataset across different sample size using TF method.
Figure 16: (a) Cross-evaluation using QE method with sample size N=16 on Math dataset. (b) Cross-evaluation using QE method with sample size N=16 on GSM8k dataset. For each solver model, sampled responses are verified by different models.
Figure 17: (a) Cross-evaluation using TF method with sample size N=16 on Math dataset. (b) Cross-evaluation using TF method with sample size N=16 on GSM8k dataset. (c) Cross-evaluation using TF method with sample size N=16 on ProntoQA dataset. For each solver model, sampled responses are verified by different models. Figures on the left illustrate the difference of accuracy while figures on the right show the 10 logarithms of the uncertainty difference.
(a) TF
(b) QE
Figure 18: Pass@K with TF and QE method for different $K$ at $t=0$, $t=5$ and $t=10$. Pass@K evaluates the diversity of model generation. The efficacy of the self-improvement process is demonstrated by an increase in the Pass@K metric for small values of $K$. Conversely, for large values of $K$, a decrease in Pass@K is observed. This phenomenon suggests that the diversity of generations decreases during the self-improvement process.